OpenStudy (msmr):

Determine the function f(x), where f(x) is a quadratic function with zeros f(2)=f(-4)=0 and y intercept of 6.

OpenStudy (anonymous):

start with \(f(x)=a(x-2)(x+4)\) because that will have the appropriate zeros then solve for \(a\) by saying \(f(0)=6=a(-2)(4)\)

OpenStudy (msmr):

why does that function have the appropriate zeros? I don't really understand this...

OpenStudy (anonymous):

ok you know if you want to find the zeros of some quadratic the first step would be to try and factor right?

OpenStudy (msmr):

right, i understand that

OpenStudy (msmr):

like if the zeros of a function were 3 and 2, wouldn't f(x) = (x-3)(x+2)?

OpenStudy (anonymous):

no it would be \((x-3)(x-2)\)

OpenStudy (anonymous):

because if you replace \(x\) by 3 you get zero for the first factor, and if you replace x by 2 you get zero for the second factor

OpenStudy (msmr):

i meant to say x-2, sorry, i mistyped that. that part i do understand.

OpenStudy (anonymous):

so this is the same idea, but in reverse instead of factoring to find the zeros, you are given the zeros and so you know how to factor

OpenStudy (anonymous):

it is called the "factor theorem" if \(r\) is a zero of a polynomial \(p(x)\) then you can factor as \[p(x)=(x-r)q(x)\] for some polynomial \(q\) in other words it works both ways factor, find the zeros have the zeros, factor

OpenStudy (msmr):

but why does it say f(2)=f(-4)? how can f(2) equal f(-4)?

OpenStudy (anonymous):

oh i think you are confused by the wording of the question you are TOLD that \(f(2)=0\) and that \(f(-4)=0\)

OpenStudy (anonymous):

so they could have said "the zeros are at \(2\) and \(-4\)" or "the \(x\) intercepts are at \(2\) and \(-4\)" or "\(f(2)=0\) and also \(f(-4)=0\)" but to be succinct they just wrote \[f(2)=f(-4)=0\]

OpenStudy (msmr):

I'm so sorry I disappeared, my internet went down and I had to restart my computer to fix it.

OpenStudy (msmr):

I was confused byt he wording, thank you. If you're still there, I have another question...using your equation, I got a = -3/4 The answer provided to me says y = -3/4x^2 - 3/2x + 6

OpenStudy (msmr):

and I am a bit confused on how they got the 3/2 part

OpenStudy (anonymous):

i think they multiplied out

OpenStudy (anonymous):

you find \(a(-2)(4)=6\) so \(a=-\frac{3}{4}\) which looks good to me

OpenStudy (anonymous):

then \[-\frac{3}{4}(x-2)(x+4)=-\frac{3}{4}(x^2+2x-8)=-\frac{3}{4}x^2-\frac{3}{2}x+6\]

OpenStudy (msmr):

thank you! that was very helpful. again, i'm sorry i abandoned you for a while!

OpenStudy (anonymous):

np, i went away too